VEGAS Tests 2014 Jan 02

Goals

  • Test modes 1-29 with stationary telescope and LO1B test tones, astronomical test observations for AAS and for modes 20-29

Observing details

  • session number is 26
  • Bank A and B are in a fault state. Try restarting them through Device Explorer; no effect. Can turn them off and on, this clears out the fault status
  • This clears the Managers, but scan 1 fails with: "No data received from vegas_hp_server - call Joe. He reinstalls and restarts the VEGAS HPC server
  • Scan 2 - Lbandmode20TestTone same problem
  • Scan 3 - Track - same problem - Joe thinks this is a GPU resource problem and reinitializes them
  • Scan 4 - Track - aborts: "deadline missed, arm time is in the past"
  • Joe runs some scans from Device Explorer, then reboots HPC machines
  • Scan 5 - Track - seems to work ok (didn't reconfigure - this configuration is what was set by Joe - 1 sec exposure x 10 seconds)
  • Scan 6 - Lbandmode20ToneTest - appears to work, but config resets LO1B - only one integration
  • Scan 7 Track forget to set Tone B
  • Scan 8 Track. Fill and plot with GBTIDL - need to remember plnum - all looks correct!
  • Scan 9, 10 Mode 20, using setValues to set the LO options
  • Scan 11 - Mode 21 - doesn't seem to have some data in some IFNUMs, only seems to have one integration - as expected.I may have just run GBTIDL
  • Change tint = 2.0, scanlength = 20
  • Scan 12 - Mode 20 - Bank B is fine, but Bank A aborts - "no data received from vegas_hpc_server" got 4 integrations for Bank B, but the IFNUMs are not in the order I would expect. Ignore this scan as pathological. Joe investigates why Bank A aborted. Finds an error
  • See the test tone, and a host of other lines, try setting tone down to -20dB
  • Scan 13 - Mode 20 - looks good, data in all IFNUMs
  • Scan 14 - Mode 21 - looks good also
  • Scan 15 - Mode 22 - looks good
  • Scan 16 - Mode 23 - looks good
  • Scan 17 - Mode 24 - all cal='T' data seems to have been blanked for the first integration only? Probably due to dropped packets. Otherwise ok.
  • Scan 18 - Mode 25 - looks good
  • Scan 19 - Mode 26 - looks good
  • Scan 20 - Mode 27 - looks good
  • Scan 21 - Mode 28 - looks good
  • Scan 22 - Mode 29 - only 9 integrations, otherwise looks good. 9 integrations is correct according to Joe, based on actual exposure time.
  • Switch to using LbandssmodeToneTest (ss = single subband). Edit as appropriate to switch through modes.
  • Scan 23 - Mode 1 - need to move tone
  • Scan 24 - Mode 2 - need to move tone
  • Joe runs some scans over lunch - not sure of the scan numbers
  • Scan 25 - mode 1 - don't see a tone at 1460 MHz - maybe wider bandwidth means this is less obvious? Try turning up the tone...
  • Scan 26 - mode 1 - now seems present at 1450 MHz - try turning it off, to confirm that it is -
  • Scan 27 - mode 1 - yes, the line goes away.
  • Joe realizes that sub_frequencyA, etc are set to 760000000 (see in dev explorer) but should be 750000000. Put the "!setValues" line back in.
  • Scan 28 - mode 1 - sub_frequencyA does not seem to have been set. Looks like there were some lines missing... Try and recreate them.
  • Scan 29 - mode 1 - line now appears at correct place, 1460 MHz, no strong ripples, just the passband
  • Scan 30 - mode 2 - ok - no sign of strong ripples, just the passband
  • Scan 31 - mode 3 - ok - no sign of strong ripples, just the passband
  • Scan 32 - mode 4 - ok, but strong ripples (six across the band)
  • Scan 33 - mode 5 - ok, but strong ripples
  • Scan 34 - mode 6 - ok
  • Scan 35 - mode 5 again by mistake
  • Scan 36 - mode 7 - ok, strong ripples
  • Scan 37 - mode 8 - only one IF - HPC manager crashed?
  • Scan 38 - repeat mode 8 - ok, strong ripples
  • Scan 39 - mode 9 - ok, strong ripples
  • Scan 40 - mode 10 - incorrect frequency scale, but in this case I should not have been setting sub_frequency
  • Scan 41 - mode 10 - correct frequency scale now. Line at 1425, but also a strong line at 1413.286 - the radar line?
  • Scan 42 - mode 10. Test tone moves to 1424, but other line moves to 1412.28 - so what could this be? Three strong ripples
  • Scan 43 - mode 11 - ok, three strong ripples, mystery line
  • Scan 44 - mode 12- ok, as mode 11
  • Scan 45 - mode 13 - ok, as mode 11
  • Scan 46 - mode 14 - ok, as mode 11
  • Scan 47 - mode 15 - ok, as mode 11
  • Scan 48 - mode 16 - ok, as mode 11
  • Scan 49 - mode 17 - ok, as mode 11
  • Scan 50 - mode 18 - ok, as mode 11
  • Scan 51 - mode 19 - ok, as mode 11 !!!
  • Scan 52 - mode 19, no test tone, no spurious lines apart from the bandwidth/2 ADC effect
  • Scan 53 - mode 20 frequency switched, 0,5 MHz switch offset, Tint = 2.0. I think this all looks correct
  • Scan 54 - mode 20 Tp with CAL, 10ms switch period, 2 second integration
  • Scan 55 - mode 20 TP with CAL, 10 ms switch, 10 ms integration, manager rounded this up to ~20ms.

Astronomical Observations

  • switch to session 27
  • snow in the dish, so observe at L-band initially; still some light snow.
  • Scan 1/2 Lbandmode20OnOff - galaxy F81
  • Scans 3-12 repeat
  • switch to X-band
  • switch to mode 2 - won't configure. Need to restart turtle, and turn managers off and on
  • Scan 13-17 peak/focus on 0115-0127
  • Scan 18/19 on-off on 3C48 - looks good
  • Scan 20-24 - p/f on 0207+6246
  • Scan 25-34 on-offs on W3 (Lbandmode2OnOff)
  • Scan 35-36 on-off on W3 (Lbandmode2OnOff) with frequency offset to 9500 MHz
  • Try Amanda's mode 29 script - won't configure. Maybe too much total bandwidth requested? Amanda fixes it
  • Scan 37-46 mode 29
  • no obvious signal - switch back to L-band
  • Scan 47-50 peak at L-band
  • 51-60 OnOff using Lbandmode20OnOff)
  • switch to another HI Galaxy - U2855
  • 61-64 - peak/focus
  • 65-74 - OnOffs - something is really mixed up in Scans 65/66 <=== need to check this with Bob
  • pack in at 7:10pm

Topic revision: r4 - 2014-01-03, RichardPrestage
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback